hi again..
1. A hospital needs to dilute a 50% boric acid solution to a 10% solution. If it needs 25 liters of the 10% solution, how much of the 50% solution and how much water should it use?
2. anabel has 25 ounces of a boric acid solution which she wishes to dilute to a 10% solution. how much of the 10% solution must be used?

Answer by stanbon(75887)

1. A hospital needs to dilute a 50% boric acid solution to a 10% solution. If it needs 25 liters of the 10% solution, how much of the 50% solution and how much water should it use?
-----------------------
Equation:
boric + boric = boric
0.50x + 0*(25-x) = 0.10*25
---
Multiply thru by 100 to get:
50x = 10*25
x = 5 liters (amt. of 50% solution needed for the final mixture)
25-x = 20 liters (amt of water needed for the final mixture)
=====================
2. anabel has 25 ounces of a boric acid solution which she wishes to dilute to a 10% solution. how much of the 10% solution must be used?
------
If you start with 100% boric acid and then add a 10% solution you
cannot end up with a 10% solution.
--------------------------
Cheers,
Stan H.